OpenStudy (anonymous):

I just can't seem to get this right help? Simplify the following expression. 5x - 6y + 5z - x - 12y + 2z

OpenStudy (anonymous):

do you know what like terms are ?

OpenStudy (anonymous):

yes but i must be confusing myself somewhere

OpenStudy (anonymous):

like terms are, well, terms that have the same variable. For instance, 6y and 12y are like terms. You cannot mix terms that are not alike. You cannot add 5x + 5z because the terms are not alike. So lets get our like terms together... 5x - x = 4x -6y - 12y = - 18y 5z + 2z = 7z now put them all together... 4x - 18y + 7z